OpenStudy (anonymous):

Kim has $1.65 in nickels, dimes and quarters. She has 10 coins all together and the number of quarters is equal to the number of nickels and dimes combined. Haw many of each coin does she have?

OpenStudy (anonymous):

let x be the number of nickels, let y be the number of dimes, since same number of quarters as nickels, x is also the number of quarters

OpenStudy (anonymous):

x + y + x =10 total number of coins 0.05x + 0.1y+0.25x=1.65 total value of coins

OpenStudy (anonymous):

solving these two equations gives you the number of each coin

OpenStudy (anonymous):

idk how to solve it dude.

OpenStudy (anonymous):

the first equation is 2x + y = 10 the second equation is 0.3x+0.1y=1.65